OpenStudy (anonymous):

Find the solution set of this inequality.Enter your answer in interval notation using grouping symbols. |-2x+1|<=5

OpenStudy (anonymous):

first convince yourself that |-2x+1|=|2x-1| this will make life easier. then write \[|2x-1|\leq 5\] \[-5\leq 2x-1\leq 5\] \[-4\leq 2x\leq6\] \[-2\leq x\leq 3\]

OpenStudy (anonymous):

you will get the same answer just starting with \[-5\leq -2x+1\leq 5\] but it will be more annoying because of the -2 in front of the x. try it and see what happens

OpenStudy (anonymous):

good morning senesei. how is the toyota?

OpenStudy (anonymous):

i understand it now thanks